# Auslander-Reiten translations in the monomorphism categories of exact categories

### v1

@@ 2.1. Exact categories and (co)stable categories
We are mostly interested in additive categories which are Krull-Schmidt. Recall that an additive category $\mathcal{C}$ is called Krull-Schmidt if any object $M\in\mathcal{C}$ has a unique (up to permutation) decomposition: $M=\mathop{\oplus}\limits_{i=1}^{n}M_{i}$ where $\operatorname{End}_{\mathcal{C}}(M_{i})$ is a local ring for all $1\leq i\leq n$.
The following result is well known (e.g. [KS, Proposition 1.2]), which guarantees the existence of projective covers and injective envelopes in Krull-Schmidt categories.
@@
Let ${\mathcal{C}}$ be a Krull-Schmidt exact category with enough projective (resp. injective) objects. Then each object admits a projective cover (resp. injective envelope).
Denote by ${\mathcal{C}}_{\mathcal{P}}$ (resp. ${\mathcal{C}}_{\mathcal{I}}$) the full subcategory consisting of objects in ${\mathcal{C}}$ which do not contain any non-zero projective (resp. injective) summands. There is an equivalence $\underline{{\mathcal{C}}_{\mathcal{P}}}\cong\underline{{\mathcal{C}}}$ (resp. $\overline{{\mathcal{C}}_{\mathcal{I}}}\cong\overline{\mathcal{C}}$) induced by the inclusion ${\mathcal{C}}_{\mathcal{P}}\to{\mathcal{C}}$ (resp. ${\mathcal{C}}_{\mathcal{I}}\to{\mathcal{C}}$). The following result is easy to prove (cf.[ARS, IV. Exercise 7]).
@@
Let ${\mathcal{C}}$ be a Krull-Schmidt exact category with enough projectives. A morphism $f:A\to B$ in ${\mathcal{C}}_{\mathcal{P}}$ is an isomorphism if and only if $\underline{f}$ is an isomorphism in $\underline{{\mathcal{C}}}$.
@@
Let $\iota$ be the inclusion of ${\mathcal{C}}$ as a full subcategory of ${\mathcal{D}}$. The restriction $Q_{{\mathcal{C}}}|_{{\mathcal{C}}_{{\mathcal{P}}}}=Q_{\mathcal{C}}\circ\iota$ is an epivalence.
@@ 2.2. Almost split sequences
Following [INP, Definition 3.4], we introduce the Auslander-Reiten-Serre duality for exact categories. An additive category ${\mathcal{C}}$ is $k$-linear if its hom-sets are $k$-modules. Denote by $D=\operatorname{Hom}_{k}(-,k)$ the duality functor.
@@
(1) A right Auslander-Reiten-Serre (ARS) duality is a pair $(\tau,\eta)$ of an additive functor $\tau:\underline{{\mathcal{C}}}\to\overline{\mathcal{C}}$ and a binatural isomorphism $\eta_{A,B}:\operatorname{Hom}_{\underline{{\mathcal{C}}}}(A,B)\cong D\operatorname{Ext}^{1}_{\mathcal{C}}(B,\tau A)$ for any $A,B\in{\mathcal{C}}$
@@
Let ${\mathcal{C}}$ be a Krull-Schmidt exact category. We call $f:A\to B$ in ${\mathcal{C}}$ a source map or left almost split, if it is not a section and each $u:A\to B^{\prime}$ which is not a section factors through $f$. Dually, we call $g:B\to C$ a sink map or right almost split, if it is not a retraction and each $v:B^{\prime}\to C$ which is not a retraction factors through $g$.
An almost split sequence in ${\mathcal{C}}$ is a conflation $A\stackrel{{\scriptstyle f}}{{\to}}B\stackrel{{\scriptstyle g}}{{\to}}C$ such that $f$ is a source map and $g$ is a sink map.
We say ${\mathcal{C}}$ has almost split sequences if it satisfies:
@@
    Each indecomposable object $A$ admits a source map $A\to B$ and a sink map $B\to A$.
@@
    For each indecomposable non-projective object $C$, there is an almost split sequence $A\to B\to C$.

### v2

@@ 2.1. Exact categories and (co)stable categories
2.2. Krull-Schmidt categories.
We are mostly interested in additive categories which are Krull-Schmidt. Recall that an additive category $\mathcal{C}$ is called Krull-Schmidt if any object $M\in\mathcal{C}$ has a unique (up to permutation) decomposition: $M=\mathop{\oplus}\limits_{i=1}^{n}M_{i}$,where $\operatorname{End}_{\mathcal{C}}(M_{i})$ is a local ring for all $1\leq i\leq n$. It is well-known that a category is Krull-Schmidt if and only if it has split idempotents and the endomorphism ring of each object is semi-perfect [Kr, Corollary 4.4]. In particular, $\operatorname{Hom}$-finite idempotent split $k$-linear categories are Krull-Schmidt.
Recall that in an additive category $\mathcal{C}$}, two morphisms $f:X\to Y$ and $f^{\prime}:X^{\prime}\to Y$ are called right equivalent if there exist $g:X\to X^{\prime}$ and $h:X^{\prime}\to X$ such that $f=f^{\prime}g$ and $f^{\prime}=fh$. A morphism $f:X\to Y$ is called right minimal if any endomorphism $g\in\operatorname{End}(X)$ such that $f=fg$ is an automorphism. It is well-known that in a Krull-Schmidt category, any morphism $f:X\to Y$ has a unique up to right equivalence decomposition $f=(f_{1},f_{2}):X=X_{1}\oplus X_{2}\to Y$ such that $f_{1}$ is right minimal and $f_{2}=0$ (see [KS, Corollary 1.4]), where $f_{1}$ is called the right minimal version of $f$. The notion of left minimal mophisms is defined dually.
A morphism $f:P\to M$ in an exact category $\mathcal{C}$ is called an $\operatorname{Ext}$-projective cover, if $f$ is a right minimal deflation and $P$ is $\operatorname{Ext}$-projective. The dual notion of $\operatorname{Ext}$-injective envelopes are defined similarly. Due to the existence of right (left) minimal morphisms in Krull-Schmidt categories, it is easy to see the existence of $\operatorname{Ext}$-projective cover and $\operatorname{Ext}$-injective envelopes.
@@
Let $\mathcal{C}$} be a Krull-Schmidt exact category with enough $\operatorname{Ext}$-projective (resp. $\operatorname{Ext}$-injective) objects. Then each object admits an $\operatorname{Ext}$-projective cover (resp. $\operatorname{Ext}$-injective envelope).
Denote by ${\mathcal{C}}_{\mathcal{P}}$ (resp. $\mathcal{C}$$_{\mathcal{I}}$) the full subcategory consisting of objects in $\mathcal{C}$} which do not contain any non-zero $\operatorname{Ext}$-projective (resp. $\operatorname{Ext}$-injective) summands. There is an equivalence $\underline{{\mathcal{C}}_{\mathcal{P}}}\cong\underline{{\mathcal{C}}}$ (resp. $\overline{{\mathcal{C}}_{\mathcal{I}}}\cong\overline{{\mathcal{C}}}$) induced by the inclusion ${\mathcal{C}}_{\mathcal{P}}\to{\mathcal{C}}$ (resp. ${\mathcal{C}}_{\mathcal{I}}\to{\mathcal{C}}$). The following result is easy to prove (cf.[ARS, IV. Exercise 7]).
@@
Let $\mathcal{C}$} be a Krull-Schmidt exact category with enough $\operatorname{Ext}$-projectives. A morphism $f:A\to B$ in ${\mathcal{C}}_{\mathcal{P}}$ is an isomorphism if and only if $\underline{f}$ is an isomorphism in $\underline{{\mathcal{C}}}$.
@@
Let $\mathcal{C}$} be a Krull-Schmidt exact category with enough $\operatorname{Ext}$-projectives. Let $\iota$ be the inclusion of ${\mathcal{C}}_{\mathcal{P}}$ as a full subcategory of $\mathcal{C}$}. The restriction $Q_{{\mathcal{C}}}|_{{\mathcal{C}}_{{\mathcal{P}}}}=Q_{\mathcal{C}}\circ\iota:{\mathcal{C}}_{\mathcal{P}}\to\underline{{\mathcal{C}}}$ is an epivalence.
@@ 2.3. Almost split sequences
Following [INP, Definition 3.4], we introduce the Auslander-Reiten-Serre duality for exact categories.
@@
(1) A right Auslander-Reiten-Serre (ARS) duality is a pair $(\tau,\eta)$ of an additive functor $\tau:\underline{{\mathcal{C}}}\to\overline{{\mathcal{C}}}$ and a binatural isomorphism $\eta_{A,B}:\operatorname{Hom}_{\underline{{\mathcal{C}}}}(A,B)\cong D\operatorname{Ext}^{1}_{\mathcal{C}}(B,\tau A)$ for any $A,B\in{\mathcal{C}}$ .
@@
Let $\mathcal{C}$} be an additive category. We call $f:A\to B$ in $\mathcal{C}$} a source map or left almost split, if it is not a section and each $u:A\to B^{\prime}$ which is not a section factors through $f$. Dually, we call $g:B\to C$ a sink map or right almost split, if it is not a retraction and each $v:B^{\prime}\to C$ which is not a retraction factors through $g$. Furthermore, if $\mathcal{C}$} is an exact category. An almost split sequence is a conflation $A\stackrel{{\scriptstyle f}}{{\to}}B\stackrel{{\scriptstyle g}}{{\to}}C$ such that $f$ is minimal left almost split and $g$ is minimal right almost split. We say an exact category $\mathcal{C}$} has left (right) almost split sequences if every endo-local non-$\operatorname{Ext}$-injective (projective) object is the starting (ending) term of an almost split sequence.
Notice that the definitions of source maps and sink maps are based on the homomorphisms of the category $\mathcal{C}$} and irrelevant to the exact structures over $\mathcal{C}$}. However, we want to emphasize that almost split sequences depend not only on the category but also on the exact structure.
Lemma 2.12.
Let $\mathcal{E}$$\subseteq{\mathcal{E}}^{\prime}$ be two exact structures on $\mathcal{C}$}. If $\delta:0\to A\to B\to C\to 0$ is an almost split sequence in $\mathcal{E}$}, then it is still almost split in $\mathcal{E}$$^{\prime}$.
3. Morphism categories and functor categories
3.1. Exact structures on morphism categories
Definition 3.1 (The deg-exact structure).
Let $({\mathcal{C}},{\mathcal{E}})$ be an exact category. Let ${\mathcal{E}}^{\rm deg}$ be the class of all pairs of composable morphisms
${{\left(\begin{smallmatrix}X_{1}\\ X_{2}\end{smallmatrix}\right)}_{f}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\left(\begin{smallmatrix}\phi_{1}\\ \phi_{2}\end{smallmatrix}\right)}$${{\left(\begin{smallmatrix}Y_{1}\\ Y_{2}\end{smallmatrix}\right)}_{g}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\left(\begin{smallmatrix}\psi_{1}\\ \psi_{2}\end{smallmatrix}\right)}$${{\left(\begin{smallmatrix}Z_{1}\\ Z_{2}\end{smallmatrix}\right)}_{h}}$
in ${\mathcal{H}}({\mathcal{C}})$ such that the induced composable morphisms $X_{i}\stackrel{{\scriptstyle\phi_{i}}}{{\rightarrow}}Y_{i}\stackrel{{\scriptstyle\psi_{i}}}{{\rightarrow}}Z_{i}$ are in $\mathcal{E}$} for $i=1,2$.
It is straightforward to see that ${\mathcal{E}}^{\rm deg}$ defines an exact structure on ${\mathcal{H}}({\mathcal{C}})$, which is usually referred to as the degreewise exact structure. In particular, for any additive category $\mathcal{C}$} there is an exact structure ${\mathcal{E}}^{\rm split}$ as mentioned in Example 2.1. Then the exact category $({\mathcal{C}},{\mathcal{E}}^{\rm split})$ induces an exact structure on ${\mathcal{H}}({\mathcal{C}})$ which can be explicitly described as below:
Definition 3.2 (The cw-exact structure).
Let ${\mathcal{E}}^{\rm cw}$ be the class of all pairs of composable morphisms
such that the induced composable morphisms $X_{i}\stackrel{{\scriptstyle\phi_{i}}}{{\rightarrow}}Y_{i}\stackrel{{\scriptstyle\psi_{i}}}{{\rightarrow}}Z_{i}$ split in $\mathcal{C}$} for $i=1,2$. The elements in ${\mathcal{E}}^{\rm cw}$ are called cw-exact sequences.
From the definition, cw-exact structure arises naturally as a $\rm deg$-exact structure.
Proposition 3.3.
Let $\mathcal{C}$} be an additive category. Then there is an equivalence of exact categories $({\mathcal{H}}({\mathcal{C}}),{\mathcal{E}}^{\rm cw})\cong({\mathcal{H}}({\mathcal{C}}^{split}),{\mathcal{E}}^{\rm deg})$.
Remark 3.4.
Any cw-exact sequence
is isomorphism to
When $\mathcal{C}$} is an exact category, there are two exact structures on ${\mathcal{H}}({\mathcal{C}})$, namely ${\mathcal{E}}^{\rm cw}$ and ${\mathcal{E}}^{\rm deg}$. We denote the exact category $({\mathcal{H}}({\mathcal{C}}),{\mathcal{E}}^{\rm deg})$ and $({\mathcal{H}}({\mathcal{C}}),{\mathcal{E}}^{\rm cw})$ by ${{\mathcal{H}}}({\mathcal{C}})^{\rm deg}$ and ${{\mathcal{H}}}({\mathcal{C}})^{\rm cw}$ respectively. It is clear that ${\mathcal{E}}^{\rm cw}\subseteq{\mathcal{E}}^{\rm deg}$. So any cw-exact sequence is still exact in $({\mathcal{H}}({\mathcal{C}}),{\mathcal{E}}^{\rm deg})$. Intuitively, ${\mathcal{E}}^{\rm deg}$ is the more natural exact structure because when ${\mathcal{C}}\subseteq\Lambda\textendash\operatorname{mod}$ is an exact subcategory, $({\mathcal{H}}({\mathcal{C}}),{\mathcal{E}}^{\rm deg})$ is an exact subcategory of ${\mathcal{H}}(\Lambda)\cong T_{2}(\Lambda)\textendash\operatorname{mod}$. Hence, in the following, when we refer to the exact category ${{\mathcal{H}}}({\mathcal{C}})$ without emphasizing the exact structure, we mean $({\mathcal{H}}({\mathcal{C}}),{\mathcal{E}}^{\rm deg})$. Although the exact structure ${\mathcal{E}}^{\rm cw}$ is much smaller than ${\mathcal{E}}^{\rm deg}$ in general, it is surprising that when $({\mathcal{H}}({\mathcal{C}}),{\mathcal{E}}^{\rm deg})$ has almost split sequences, most of the almost split sequences are cw-exact. Hence, when it comes to characterizing almost split sequences, ${\mathcal{H}}({\mathcal{C}})^{\rm cw}$ becomes a convenient tool.
3.2. Functors on morphisms categories
Proposition 3.5.
(1) Let $F:{\mathcal{C}}\to{\mathcal{D}}$ be an additive functor. If $F$ is an equivalence, then so is ${\mathcal{H}}(F)$. (2) Let $F:{\mathcal{C}}\to{\mathcal{D}}$ be an additive functor between exact categories. Then $F$ preserves (reflects) exact sequences if and only if ${\mathcal{H}}(F)$ preserves (reflects) exact sequences.
Proof.
(1) It suffices to prove ${\mathcal{H}}(F)$ is dense and fully-faithful. For the density of ${\mathcal{H}}(F)$, let $\left(\begin{smallmatrix}A^{\prime}_{1}\\ A^{\prime}_{2}\end{smallmatrix}\right)_{g}\in{\mathcal{H}}({\mathcal{D}})$. By the density of $F$, there are $A_{1},A_{2}\in{\mathcal{C}}$, such that $F(A_{1})\cong A^{\prime}_{1},F(A_{2})\cong A^{\prime}_{2}$. Since $F$ is full, there is a homomorphism $f:A_{1}\to A_{2}$ such that $F(f)\cong g.$ So ${\mathcal{H}}(F)$ is dense.
For the fullness of ${\mathcal{H}}(F)$, let $\left(\begin{smallmatrix}A_{1}\\ A_{2}\end{smallmatrix}\right)_{f}$ and $\left(\begin{smallmatrix}B_{1}\\ B_{2}\end{smallmatrix}\right)_{g}$ be objects in ${\mathcal{H}}({\mathcal{C}})$ and $(\alpha^{\prime},\beta^{\prime}):Ff\to Fg$ be a homomorphism in ${\mathcal{H}}({\mathcal{D}})$, i.e. $Fg\alpha^{\prime}=\beta^{\prime}Ff$. By the fullness of $F$, there are homomorphisms $\alpha:A_{1}\to B_{1}$ and $\beta:A_{2}\to B_{2}$ such that $F\alpha=\alpha^{\prime}$ and $F\beta=\beta^{\prime}$. Hence $FgF\alpha=F\beta Ff$. Since $F$ is a faithful functor, we have $g\alpha=\beta f$. That is to say, $(\alpha,\beta):f\to g$ is a homomophism in ${\mathcal{H}}({\mathcal{C}})$ such that $F(\alpha,\beta)=(\alpha^{\prime},\beta^{\prime})$. Hence ${\mathcal{H}}(F)$ is full. It is easy to see that the faithfulness of ${\mathcal{H}}(F)$ is guaranteed by the faithfulness of $F$. This completes the proof of (1). Statement (2) is straightforward by the definition of the exact structure ${\mathcal{E}}^{\rm deg}$. ∎
Let $\mathcal{C}$} and $\mathcal{D}$} be additive categories. Since any additive functor $F:{\mathcal{C}}\to{\mathcal{D}}$ preserves split exact sequences. It follows immediately from (2) that ${\mathcal{H}}(F)$ is an exact functor ${\mathcal{H}}({\mathcal{C}})^{\rm cw}\to{\mathcal{H}}({\mathcal{D}})^{\rm cw}$.
Let $\mathcal{C}$} be an exact category. Define the monomorphism category $\mathcal{S}({\mathcal{C}})$ (resp. epimorphism category $\mathcal{F}({\mathcal{C}})$) as the full subcategory of $(\mathcal{H}({\mathcal{C}}),{\mathcal{E}}^{\rm deg})$, consisting of objects which are inflations (resp. deflations) in $\mathcal{C}$}. Notice that different exact structures on $\mathcal{C}$} may result in different monomorphism (epimorphism) categories.
There are a couple of functors:
Proposition 3.6.
Let $\mathcal{C}$} be an exact category. Additive functors $\operatorname{\bf Ker}$ and $\operatorname{\bf Cok}$ form a pair of inverse equivalences.
As immediate consequences of the RSS equivalence. We see that $\operatorname{\bf Ker}$ and $\operatorname{\bf Cok}$ are exact equivalences which preserve indecomposable objects, $\operatorname{Ext}$-projectives, $\operatorname{Ext}$-injectives and almost split sequences.
As above, for an exact functor $F:{\mathcal{C}}\to{\mathcal{D}}$, we can consider the induced functors ${\mathcal{S}}(F):{\mathcal{S}}({\mathcal{C}})\to{\mathcal{S}}({\mathcal{D}})$ and ${\mathcal{F}}(F):{\mathcal{F}}({\mathcal{C}})\to{\mathcal{F}}({\mathcal{D}})$, which send $\left(\begin{smallmatrix}A\\ B\end{smallmatrix}\right)_{f}$ to $\left(\begin{smallmatrix}FA\\ FB\end{smallmatrix}\right)_{Ff}$ and send morphisms $(\alpha,\beta):f\to g$ to $(F\alpha,F\beta):Ff\to Fg$. It is easy to see that ${\mathcal{S}}(F)$ and ${\mathcal{F}}(F)$ are also exact.
Proposition 3.7.
Let $F:{\mathcal{C}}\to{\mathcal{D}}$ be an exact functor. If $F$ is an exact equivalence, then ${\mathcal{S}}(F)$ and ${\mathcal{F}}(F)$ are also exact equivalences.
Proof.
To show that ${\mathcal{S}}(F)$ and ${\mathcal{F}}(F)$ are fully-faithful is similar to the proof of Proposition 3.5, which we omit. To see ${\mathcal{S}}(F)$ (or ${\mathcal{F}}(F))$ is dense, notice that since $F$ is full and dense, for any object $\left(\begin{smallmatrix}A\\ B\end{smallmatrix}\right)_{f}$ in ${\mathcal{S}}({\mathcal{D}})$ (resp. ${\mathcal{F}}({\mathcal{D}})$), there is an object $\left(\begin{smallmatrix}X\\ Y\end{smallmatrix}\right)_{u}$ in ${\mathcal{H}}({\mathcal{C}})$ such that $\left(\begin{smallmatrix}FX\\ FY\end{smallmatrix}\right)_{Fu}\cong\left(\begin{smallmatrix}A\\ B\end{smallmatrix}\right)_{f}$. But $u$ must be a inflation (resp. deflation) as $F$ reflects exact sequences. Clearly, both ${\mathcal{S}}(F)$ and ${\mathcal{F}}(F)$ preserve and reflect exact sequences. ∎
3.3. Almost split sequences in morphism categories
In the following, we will characterize the almost split sequences in morphism categories.
There are some special almost split sequences in the morphism categories which are easy to check by definitions. We call them the sporadic almost split sequences.
Lemma 3.8.
Let $\mathcal{C}$} be an exact category. For objects $A,C\in{\mathcal{C}}$, let $p:P(C)\to C$ be the $\operatorname{Ext}$-projective cover and $e:A\to I(A)$ be the $\operatorname{Ext}$-injective envelope. Assume $\delta$ is an almost split sequence in $\mathcal{C}$}, $p^{\prime}$ is a lift of the $\operatorname{Ext}$-projective cover $p$ and $e^{\prime}$ is a lift of the $\operatorname{Ext}$-injective envelope $e$.
[…]
@@ 3.3. Almost split sequences in morphism categories
    [HE, Lemma 4.5] The following exact sequences are almost split in ${{\mathcal{H}}}({\mathcal{C}})$.
    ${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\left(\begin{smallmatrix}A\\ A\end{smallmatrix}\right)}_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\left(\begin{smallmatrix}1\\ f\end{smallmatrix}\right)}$${{\left(\begin{smallmatrix}A\\ B\end{smallmatrix}\right)}_{f}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\left(\begin{smallmatrix}0\\ g\end{smallmatrix}\right)}$${{\left(\begin{smallmatrix}0\\ C\end{smallmatrix}\right)}_{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$
    ${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\left(\begin{smallmatrix}A\\ 0\end{smallmatrix}\right)}_{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\left(\begin{smallmatrix}f\\ 0\end{smallmatrix}\right)}$${{\left(\begin{smallmatrix}B\\ C\end{smallmatrix}\right)}_{g}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\left(\begin{smallmatrix}g\\ 1\end{smallmatrix}\right)}$${{\left(\begin{smallmatrix}C\\ C\end{smallmatrix}\right)}_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$
@@
    The following exact sequences are almost split in ${{\mathcal{S}}}({\mathcal{C}})$.
    ${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\left(\begin{smallmatrix}A\\ A\end{smallmatrix}\right)}_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\left(\begin{smallmatrix}1\\ f\end{smallmatrix}\right)}$${{\left(\begin{smallmatrix}A\\ B\end{smallmatrix}\right)}_{f}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\left(\begin{smallmatrix}0\\ g\end{smallmatrix}\right)}$${{\left(\begin{smallmatrix}0\\ C\end{smallmatrix}\right)}_{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$
    ${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\left(\begin{smallmatrix}A\\ I(A)\end{smallmatrix}\right)}_{e}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\left(\begin{smallmatrix}f\\ {\left(\begin{smallmatrix}1\\ 0\end{smallmatrix}\right)}\end{smallmatrix}\right)}$${{\left(\begin{smallmatrix}B\\ I(A)\oplus C\end{smallmatrix}\right)}_{\left(\begin{smallmatrix}e^{\prime}\\ \end{smallmatrix}\right)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\left(\begin{smallmatrix}g\\ {\left(\begin{smallmatrix}1&0\end{smallmatrix}\right)}\end{smallmatrix}\right)}$${{\left(\begin{smallmatrix}C\\ C\end{smallmatrix}\right)}_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$